OpenStudy (anonymous):

What is the slope of the line that passes through the given points? (2,12), (6,11)

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

My answer is 4

OpenStudy (anonymous):

Is it 1/-4 (-4)?

pooja195 (pooja195):

Not quite... \[\LARGE \frac{11-12}{6-2}=\frac{ -1 }{ 4 } =? \]

OpenStudy (anonymous):

-4

pooja195 (pooja195):

No you leave it as -1/4

OpenStudy (anonymous):

ohhh okay thank you, could you help me with one more?

pooja195 (pooja195):

yes :)

OpenStudy (anonymous):

(6,-1),(-3,-1)

pooja195 (pooja195):

Same thing

OpenStudy (anonymous):

\[\frac{ -1-(-1) }{ 6-(-3) }\]

OpenStudy (anonymous):

i got 0/9, does that mean its undefined?

OpenStudy (anonymous):

or 0?

pooja195 (pooja195):

i think it would just be 0

OpenStudy (anonymous):

right, but if it was 9/0 that would be undefined?

OpenStudy (anonymous):

Thank you!!